OpenStudy (anonymous):

which of the following equations is equivalent to y=lnx? a x=y^e b y=x^e c x=e^y d y=e^x okkkk so im really not sure my best shot would be D ??

OpenStudy (anonymous):

Raise both sides to the base of e to get rid of the natural log.

jimthompson5910 (jim_thompson5910):

y=e^x is equivalent to ln(y) = x So you are incorrect.

jimthompson5910 (jim_thompson5910):

keep in mind that ln is a log it's a log with base 'e'

OpenStudy (anonymous):

\[y = \ln x , e^(y) = e^(\ln x), e^y = x\]